OpenStudy (anonymous):

a right triangle has legs with lengths of 5 cm and 12 cm. fine the length of the hypotenuse. a. 7 cm b. 13 cm c. 17 cm d. 34 cm

OpenStudy (anonymous):

\[5^2+12^2=h^2\] solve for h

OpenStudy (anonymous):

how do i do that , can you please explain

OpenStudy (anonymous):

for right triange use the pythagorem theory a^2 + b^2 = c^2 which c is the hypotenuese and a and b are other sides

OpenStudy (anonymous):